OpenStudy (anonymous):

Umfu-fu is 5 years older than Mosi, but 3 years younger than gunther. If mosi is 13 years old now, how old will gunther be in 2 years? Explain how u got ur answer

hero (hero):

Mosi = 13 Umfu-fu = 13 + 5 = 18 Gunther = 18 + 3 = 21

hero (hero):

In two years, Gunther will be 23

OpenStudy (anonymous):

thank youuu!